# Supervised heterogeneous network estimation via survival-based Bayesian graphical models

### v1

@@ 5 Discussion
In this study, we have employed the commonly used Gaussian mixture models for heterogeneity analysis, which show satisfactory performance. It can be of interest to explore alternative robust techniques, such as the $t$-distribution mixture model or other nonparametric mixture graphical models, for studying heavy-tailed data. Our focus in this data analysis has been mainly on the identification of gene networks using expression data. There is vast potential for a more comprehensive understanding of cancer mechanisms by incorporating other omics measurements, such as mutation and DNA methylation.
Funding
The research is supported by National Natural Science Foundation of China (12071273, 12271329, 72331005); Shanghai Rising-Star Program (22QA1403500); Shanghai Science and Technology Development Funds (23JC1402100); Shanghai Research Center for Data Science and Decision Technology; the Program for Innovative Research Team of SUFE; National Institutes of Health (CA204120), and National Science Foundation (2209685).
Disclosure Statement
The authors report there are no competing interests to declare.
Supplementary Materials
Additional supporting information may be found online in the supplementary materials section at the end of the article.
@@
Supplementary Materials for “Supervised Bayesian joint graphical model for simultaneous network estimation and subgroup identification”, including all technical proofs of the main theoretical results, and additional results of simulation studies and real data analysis.
@@
The R package superviseNet that implements the proposed approach, which is available at https://github.com/mengyunwu2020/superviseNet.
@@
Network estimation results for real data.

### v2

@@ 5 Discussion
In this study, we have employed the commonly used Gaussian mixture models for heterogeneity analysis, which show satisfactory performance. It can be of interest to explore alternative robust techniques, such as the exponential family graphical models proposed by [27] or other nonparametric graphical models, for studying heavy-tailed data. For survival modeling, we employ the Normal-error AFT model as a canonical example. Our framework is general and can be extended to more flexible survival models, such as the Weibull AFT model by appropriately modifying the likelihood function, representing a valuable direction for future research. Our focus in this data analysis has been mainly on the identification of gene networks using expression data. There is vast potential for a more comprehensive understanding of cancer mechanisms by incorporating other omics measurements, such as mutation and DNA methylation.
6 Disclosure statement
The authors have the following conflicts of interest to declare.
7 Data Availability Statement
The data that support the findings in this paper are openly available in TCGA (The Cancer Genome Atlas) at https://portal.gdc.cancer.gov/projects/TCGA-SKCM.
SUPPLEMENTARY MATERIAL
@@ 7 Data Availability Statement
Supplement to “Supervised Bayesian Joint Graphical Model for Simultaneous Network Estimation and Subgroup Identification”, including a graphical illustration of the method, algorithm details, sensitivity results, definition of $W$ in Assumption 2, technical proofs, simulation settings and results, and additional real data results.
@@
The R package superviseNet implementing the proposed method.
@@
Network estimation results for real data using the proposed approach.